OpenStudy (amyna):

find derivative: f(x)=e^x arcsinx Do i use product rule? Thanks for your help in advance!

jimthompson5910 (jim_thompson5910):

yes you use the product rule let g(x) = e^x and h(x) = arcsin(x) so f(x) = g(x) * h(x) f ' (x) = g ' (x)*h(x) + g(x) * h ' (x)

OpenStudy (amyna):

Thanks so much Jim!! :)